# KeepKV: Eliminating Output Perturbation in KV Cache Compression for Efficient LLMs Inference

### v1

@@ Abstract
Efficient inference of large language models (LLMs) is hindered by an ever-growing key-value (KV) cache, making KV cache compression a critical research direction. Traditional methods selectively evict less important KV cache entries based on attention scores or position heuristics, which leads to information loss and hallucinations. Recently, merging-based strategies have been explored to retain more information by merging KV pairs that would be discarded; however, these existing approaches inevitably introduce inconsistencies in attention distributions before and after merging, causing output perturbation and degraded generation quality. To overcome this challenge, we propose KeepKV, a novel adaptive KV cache merging method designed to eliminate output perturbation while preserving performance under strict memory constraints. KeepKV introduces the Electoral Votes mechanism that records merging history and adaptively adjusts attention scores. Moreover, it further leverages a novel Zero Inference-Perturbation Merging methods, keeping attention consistency and compensating for attention loss resulting from cache merging. KeepKV successfully retains essential context information within a significantly compressed cache. Extensive experiments on various benchmarks and LLM architectures demonstrate that KeepKV substantially reduces memory usage, enhances inference throughput by more than 2×2$\times$2 × and keeps superior generation quality even with $10\%$ KV cache budgets.
@@ 1 Introduction
Transformer-based large language models (LLMs) have demonstrated remarkable capabilities across various applications [1, 2, 3, 4, 5, 6, 7]. To accelerate inference, LLMs commonly employ a key-value (KV) cache mechanism, which stores the KV embeddings of previously processed tokens to avoid redundant computations [8, 9, 10]. However, as LLMs continue to support increasingly longer context lengths, the size of the KV cache grows rapidly, becoming a major bottleneck for inference [11]. For example, in the case of LLaMA-3-70B, a batch size of 128 with an 8K context length requires up to 320GB of KV cache memory [2]. Consequently, compressing the KV cache while preserving generation quality has become a crucial challenge.
@@
KV cache eviction is an effective approach to alleviating memory pressure. Existing studies primarily focus on selectively retaining important cache entries based on attention scores and token positions [12, 13, 14, 15, 16], as well as dynamically allocating varying cache sizes across layers and heads [17, 18, 19, 20]. However, eviction is irreversible—once a KV entry is evicted, its information is permanently lost, leading to issues such as hallucinations and context inconsistency [21]. To mitigate these drawbacks, recent studies have explored merging techniques [21, 22, 23, 24], which attempt to integrate cache entries that would otherwise be discarded into the remaining cache. CaM [21] first proposed merging the evicted value states into others. Recent methods such as D2O [23] and KVMerger [24] have proposed synchronized weighted merging strategies for both keys and values, achieving certain improvements. However, KV cache merging remains in an exploratory stage, particularly regarding merge candidate selection and merging weight computation. Consequently, existing strategies vary widely and lack solid theoretical foundations. For instance, D2O computes merging weights using cosine similarity, whereas KVMerger employs Gaussian kernel weights. We observe that existing methods inevitably introduce output perturbation. Specifically, the merged KV pair’s attention score is lower than the sum of the original scores prior to merging—a phenomenon we term "Attention Sag", as shown in Figure 1. These issues collectively underline the necessity for an efficient and theoretically grounded KV cache merging strategy.
@@
Figure 1: Illustration of KeepKV vs. Existing Methods. The three middle blocks represent KV subject to eviction/merging. (a) Eviction methods permanently discard them. (b) Merging methods integrates them into retained KV, but the result is not equivalent to the full KV, causing "Attention Sag." (c) Full KV serves as the ideal baseline. (d) KeepKV uses Electoral Votes as merging records and applies ZIP-Merging to minimize output disturbance, ensuring consistency and improving performance.
@@
To address this issue, we propose KeepKV, a novel merging method that maintains merging information and adaptively adjusts weights to eliminate output perturbation. The motivation behind KeepKV is to preserve the attention distribution and output consistency before and after merging at the current step, thus preventing any disturbances to inference. To achieve this, we first analyze the limitations of existing weighted merging methods in the context of attention computation, and propose the Electoral Votes mechanism, which records the number of times each KV pair has been merged, enabling the compressed cache to function equivalently to the original. Subsequently, we introduce Zero Inference-Perturbation Merging (ZIP-Merging), which automatically adjusts weights to compensate for any losses caused by merging, maintaining attention consistency. These designs theoretically guarantee zero attention-output perturbation at the current iteration despite compression.
Figure 2: Relative Performance on a summarization task at $5\%$ compression rate. KeepKV is closest to full KV ($100\%$).
@@
To extend KeepKV to multi-step generation and practical tasks, we replace the current iteration’s attention scores in KeepKV computations with estimations based on recent historical scores. This is motivated by our empirical observation of strong locality in attention scores, also confirmed in prior studies [21, 25]. Crucially, we provide theoretical analyses guaranteeing bounded output perturbation across multiple steps, achieving a controllable upper bound. Moreover, we offer a theoretical interpretation for prevalent similarity-based candidate selection methods, incorporating it into our design. Through theoretical derivation and extensive experiments, we demonstrate that KeepKV effectively preserves attention stability and output consistency, outperforming state-of-the-art KV cache eviction and merging methods. Across various benchmark tasks, KeepKV achieves superior performance and is able to maintain strong results even with $5\%$ KV cache budgets, as shown in Figure 2. The contributions of this paper are summarized as follows:
@@
  • We propose KeepKV, a novel adaptive KV cache merging approach designed to eliminate output perturbation caused by compression. KeepKV introduces the Electoral Votes mechanism and Zero Inference-Perturbation Merging to keep attention consistency.
@@ 2 Related Work
KV cache has become a major bottleneck for efficient LLMs inference. Post-training optimization serves as a key solution due to its real-time and extensible capabilities.[26]. Existing methods fall into three categories: quantization, eviction, and merging.
KV Cache Quantization. Quantization methods convert tensor values to lower precision to reduce bit-width. KVQuant [27] applies Per-Channel Quantization for keys and Per-Token Quantization for values. MiKV [28] introduces mixed-precision KV caching, where less critical KV are stored at lower precision. Additionally, GEAR [29] leverages low-rank matrix approximation for quantization residuals to minimize quantization loss. Our KeepKVreduces the number of cached KV pairs through merging, which is orthogonal to quantization methods and can be combined for better efficiency.
KV Cache Eviction. Eviction methods only retain more important KV entries. StreamingLLM [12] and LM-infinite [30] identifies the importance of the initial k tokens for generation. H2O [13], ScissorsHand [15] and RoCo [14] recognize crucial KV based on attention scores, while SnapKV [16] utilizes attention within an observation window. Recent works explore improved budget allocation strategies. Pyramid [17, 18] allocates more cache to lower layers, whereas AdaKV [31], HeadKV [19], and DuoAttention [20] focus on inter-head differences. However, eviction causes irreversible information loss, potentially degrading generation quality.
KV Cache Merging. KV cache merging combines less important KV entries instead of permanently discarding them. DMC [22] learns when and how to merge through training, resulting in a lack of generalization and requiring additional processing. In contrast, CaM [21] adaptively merges evicted value states into others but does not merge the corresponding keys. Recently, D2O [23] selects merge candidates and assigns merging weights based on cosine similarity between key states, while KVMerger [24] introduces a variant of the Agglomerative Hierarchical Clustering algorithm to group merge candidates and computes merging weights using Gaussian Kernel Weights. However, these methods fail to maintain attention consistency before and after weighted merging, leading to output perturbation. We propose a novel merging approach designed to eliminate output perturbation, supported by theoretical analysis and extensive comparisons.
@@ 3.1 Preliminary: Inference with KV Cache
$K_{t}=[K_{t-1},k_{t}],V_{t}=[V_{t-1},v_{t}],\quad k_{t}=x_{t}W_{k},v_{t}=x_{t}W_{v},$ (2)
$A^{t}=\text{softmax}\left(\frac{q_{t}K_{t}^{T}}{\sqrt{d}}\right),q_{t}=x_{t}W_{q};\quad o_{t}=\sum_{i=1}^{t}A_{i}^{t}v_{i}=\frac{\sum_{i=1}^{t}s^{t}_{i}v_{i}}{\sum_{i=1}^{t}s^{t}_{i}},s^{t}_{i}=e^{\frac{q_{t}k_{i}}{\sqrt{d}}}$ (3)
@@ 3.2 Rethinking KV Cache Eviction and Merging
Although current methods optimize eviction and cache allocation strategies [18, 19] to minimize output impact, they cannot eliminate the perturbation in Equation 4. Previous studies have indicated that attention is not always sparse, especially in tasks requiring full context, as shown in Figure 3. Moreover, evicted KV may become important later, but irreversible eviction leads to permanent loss.
@@
Attention Sag in KV Cache Merging. Merging methods integrate less important KV into others rather than discarding them. Existing studies typically use weighted merging [22, 23, 24]; formally, merging $(k_{e},v_{e})$ into $(k_{c},v_{c})$ is expressed as:
@@
Here, $(k_{r},v_{r})$ are the merged vectors, with weights $w_{e},w_{c}$ determined by the merging method. In D2O [23], they depend on the cosine similarity between $k_{e}$ and $k_{c}$, while in KVMerger [24], they are computed using Gaussian Kernel values. The weights satisfy the normalization condition $w_{e}+w_{c}=1$. However, this widely used convex combination method also introduces output perturbations:
Theorem 3.2.
@@ 3.3 Method: KeepKV
3.3.1 Electoral Votes and ZIP-Merging
Electoral Votes. To address Attention Sag, we propose the Electoral Votes mechanism, which records the number of merges $p_{i}$ (initialized to $1$) each KV pair undergoes. A natural analogy is the Electoral College system system, where electors hold votes proportional to their state’s population rather than a uniform share. The attention score of each KV is then scaled by its votes to approximate the original multiple KV’s influence before merging. For example, if a KV pair $(k_{r},v_{r})$ has a vote count of $p_{r}=3$, it is equivalent to three identical and independent instances of $(k_{r},v_{r})$ participating in the attention computation. Formally, the outputs before ($o_{t}$) and after merging ($o^{\prime}_{t}$) are defined as follows:
$o_{t}=\frac{\sum_{i=1}^{t}p_{i}s_{i}^{t}v_{i}}{\sum_{i=1}^{t}p_{i}s_{i}^{t}},\quad o^{\prime}_{t}=\frac{\sum_{i=1,i\neq e,c}^{t}p_{i}s_{i}^{t}v_{i}+p_{r}s_{r}^{t}v_{r}}{\sum_{i=1,i\neq e,c}^{t}p_{i}s_{i}^{t}+p_{r}s_{r}^{t}},p_{r}=p_{e}+p_{c}.$ (6)
@@ 3.3.1 Electoral Votes and ZIP-Merging
$k_{r}=\frac{(w_{e}k_{e}+w_{c}k_{c})\ln{\frac{w_{e}+w_{c}}{p_{e}+p_{c}}}}{w_{e}\ln{s_{e}^{t}}+w_{c}\ln{s_{c}^{t}}},v_{r}=\frac{w_{e}v_{e}+w_{c}v_{c}}{w_{e}+w_{c}},\quad w_{e}=p_{e}s_{e}^{t},w_{c}=p_{c}s_{c}^{t}.$ (7)
Theorem 3.3.
@@
This theorem confirms that our novel merging approach can eliminate output perturbations and complete the subtask introduced at the beginning of Section 3.2. However, its applicability remains limited to the current iteration, and extending it to multi-step generation requires additional design.
3.3.2 Extending to Multi-Step Generation
EMA Attention Scores. For ZIP-Merging to be effective in real-world multi-step generation, a solid comprehension of attention score dynamics is essential. Fortunately, empirical observations show that attention scores exhibit strong locality (Figure 3 (d)), meaning a token’s attention scores evolve smoothly across adjacent steps, which is also validated by prior studies [18, 21, 25]. From this, we employ the Exponential Moving Average (EMA) [32, 33] with bias correction, a widely used technique in time-series analysis, formulated as follows:
$\hat{s^{t}}=\frac{1}{1-{\alpha}^{t}}S^{t},\quad S^{t}=\left\{\begin{array}[]{lr}\sum_{k=t-w}^{t}(1-\alpha){\alpha}^{t-k}s^{k},\quad t=L&\\ \alpha S^{t-1}+(1-\alpha)s^{t},\quad t>L&\end{array}\right.$ (8)
Note that after the prefill stage, we compute EMA scores using a recent window of length $w$ rather than the entire sequence to obtain a more accurate estimation [16, 18]. We find that this method outperforms mainstream approaches, such as cumulative attention and sliding window averaging, in predicting attention scores. Building on this, replacing all score $s_{i}^{t}$ in Equation 7 with our EMA scores $\hat{s_{i}^{t}}$ from Equation 8 successfully achieves the extension to multi-step generation. Consequently, the future output perturbation becomes estimable and controllable. We present the following theorem and lemma(proof in Appendix A.4):
Theorem 3.5.
@@ 3.3.2 Extending to Multi-Step Generation
Similarity-driven merging. Lemma 3.6 shows that output perturbation decreases as prediction error $\epsilon$ reduces, and closer merging objects result in lower perturbation. Clearly, if the merged KV pairs are identical, retaining one pair and setting its Electoral Votes to 2 introduces no error in subsequent computations. This provides a theoretical justification for prior merging strategies favoring high-similarity KV pairs [23, 24]. Following this, we merge each evicted KV pair with the retained one having the highest cosine similarity of keys, using a predefined threshold $T$ to determine whether merging should occur, avoiding the overhead of dynamic adjustments like D2O [23].
We present the workflow of KeepKV in Figure 4. Notably, KeepKV imposes no specific constraints on cache allocation or token selection. It can directly integrate with common token selection methods by designating the merging pairs based on their eviction and retention sets, and it is also compatible with various cache allocation strategies. Thus, KeepKV demonstrates strong adaptability and can be combined with a range of mainstream cache compression methods, significantly enhancing both compression capability and generation quality.

### v2

# KeepKV: Achieving Periodic Lossless KV Cache Compression for Efficient LLM Inference

@@ Abstract
Efficient inference of large language models (LLMs) is hindered by an ever-growing key-value (KV) cache, making KV cache compression a critical research direction. Traditional methods selectively evict less important KV cache entries, which leads to information loss and hallucinations. Recently, merging-based strategies have been explored to retain more information by merging KV pairs that would be discarded; however, these existing approaches inevitably introduce inconsistencies in attention distributions before and after merging, causing degraded generation quality. To overcome this challenge, we propose KeepKV, a novel adaptive KV cache merging method designed to preserve performance under strict memory constraints, achieving single-step lossless compression and providing error bounds for multi-step compression. KeepKV introduces the Electoral Votes mechanism that records merging history and adaptively adjusts attention scores. Moreover, it further leverages a novel Zero Inference-Perturbation Merging method, compensating for attention loss resulting from cache merging. Extensive experiments on various benchmarks and LLM architectures demonstrate that KeepKV substantially reduces memory usage while successfully retaining essential context information, achieving over $2\times$ inference throughput improvement and maintaining superior generation quality even with only $10\%$ KV cache budgets.
@@ 1 Introduction
Transformer-based large language models (LLMs) have demonstrated remarkable capabilities across various applications (touvron2023llama2; jiang2023mistral7b; openai2024gpt4report; wan2024llm_survey; rozière2024codellama). To accelerate inference, LLMs commonly employ a key-value (KV) cache mechanism, which stores the KV embeddings of previously processed tokens to avoid redundant computations (Vaswani2017AttentionisAllYouNeed; Dai2019TransformerXL; Rae2019CompressiveTF). However, as LLMs continue to support increasingly longer context lengths, the size of the KV cache grows rapidly, becoming a major bottleneck for inference (kwon2023PagedAttention). For example, in the case of LLaMA-3-70B, a batch size of 128 with an 8K context length requires up to 320GB of KV cache memory (grattafiori2024llama3). Consequently, compressing the KV cache while preserving generation quality has become a crucial challenge.
@@
Prior works mainly explore two approaches for KV cache compression: eviction-based and merging-based methods, both of which are inherently lossy. Eviction-based approaches selectively retain critical cache entries using heuristics like attention scores and token positions, permanently discarding less critical entries (xiao2024StreamingLLM; zhang2023h2o; reid2024RoCo; liu2024scissorhands; li2024SnapKV; yang2024pyramidinfer) and thus causing context loss and potential hallucinations (zhang2024cam). In contrast, merging-based methods aim to integrate rather than discard KV entries to retain more information. Recent representative studies, such as CaM (zhang2024cam), D2O (wan2024D2O), and KVMerger (wang2024KVMerger), have explored strategies like weighted key-value merging to mitigate context loss. Nevertheless, these methods vary widely in merge candidate selection and merging weight computation, and lack solid theoretical foundations. We observe that existing strategies inevitably induce attention inconsistencies and output perturbation. Specifically, the merged KV pair’s attention score is lower than the sum of the original scores prior to merging—a phenomenon we term ”Attention Sag” (illustrated in Figure 1). These issues underline the necessity for an efficient and theoretically grounded KV cache merging strategy.
@@
Figure 1: Illustration of KeepKV vs. Existing Methods. The three middle blocks represent KV subject to eviction/merging. (a) Eviction methods permanently discard them. (b) Merging methods integrates them into retained KV, but the result is not equivalent to the full KV, causing ”Attention Sag.” (c) Full KV serves as the ideal baseline. (d) KeepKV uses Electoral Votes as merging records and applies ZIP-Merging to minimize output disturbance, ensuring consistency and improving performance.
In this paper, we propose KeepKV, a novel KV cache merging method designed to maintain inference consistency and preserve essential contextual information. To the best of our knowledge, KeepKV is the first approach to achieve single-step lossless compression and to provide theoretical error bounds for multi-step compression. We first conduct a comprehensive theoretical analysis of existing eviction and weighted merging methods, grounded in the attention computation process, to reveal their fundamental limitations.Building on these theoretical insights, we propose a two-stage innovative design in KeepKV:
@@
First, we propose techniques that achieve lossless compression for a single step. Specifically, we introduce the Electoral Votes mechanism, which records merging history, enabling accurate reconstruction of the original KV embeddings from compressed representations. Additionally, we present the Zero Inference-Perturbation Merging (ZIP-Merging) approach, which automatically adjusts weights to compensate for any losses caused by merging, maintaining attention consistency. These designs theoretically guarantee zero output perturbation at the current iteration despite compression.
@@
Second, we extend KeepKV to multi-step generation by estimating attention scores based on historical patterns. This is motivated by our empirical observation of strong locality in attention scores, also confirmed in prior studies (dong2024qaq; zhang2024cam). Crucially, we provide theoretical analyses guaranteeing bounded output perturbation across multiple steps, thereby ensuring consistent inference quality under extended generation. Moreover, we offer a theoretical interpretation for prevalent similarity-based candidate selection methods, incorporating it into our design.
By integrating these innovations, KeepKV further enables periodic lossless KV cache compression by storing a complete KV cache externally and periodically loading compressed representations into memory, thereby maintaining inference consistency and preserving essential contextual information. Through theoretical derivation and extensive experiments, we demonstrate that KeepKV effectively preserves attention stability and output consistency, outperforming state-of-the-art KV cache eviction and merging methods. The contributions of this paper are summarized as follows:
@@
  • We propose KeepKV, a novel adaptive KV cache merging approach, which introduces the Electoral Votes mechanism and Zero Inference-Perturbation Merging, achieving single-step lossless compression and providing error bounds for multi-step compression.
@@ 2 Related Work
KV cache has become a major bottleneck for efficient LLMs inference. Post-training optimization serves as a key solution due to its real-time and extensible capabilities.(shi2024surveyonkvcache). Existing methods fall into three categories: quantization, eviction, and merging.
KV Cache Quantization. Quantization methods convert tensor values to lower precision to reduce bit-width. KVQuant (hooper2024KVQuant) applies Per-Channel Quantization for keys and Per-Token Quantization for values. MiKV (yang2024MiKV) introduces mixed-precision KV caching, where less critical KV are stored at lower precision. Additionally, GEAR (kang2024GEAR) leverages low-rank matrix approximation for quantization residuals to minimize quantization loss. Our KeepKVreduces the number of cached KV pairs through merging, which is orthogonal to quantization methods and can be combined for better efficiency.
KV Cache Eviction. Eviction methods only retain more important KV entries. StreamingLLM (xiao2024StreamingLLM) and LM-infinite (han2024lminfinite) identifies the importance of the initial k tokens for generation. H2O (zhang2023h2o), ScissorsHand (liu2024scissorhands) and RoCo (reid2024RoCo) recognize crucial KV based on attention scores, while SnapKV (li2024SnapKV) utilizes attention within an observation window. Recent works explore improved budget allocation strategies across layers and heads. Pyramid (cai2024pyramidkv; yang2024pyramidinfer) allocates more cache to lower layers, whereas AdaKV (feng2024adakv), HeadKV (fu2024headkv), and DuoAttention (xiao2024duoattention) focus on inter-head differences. However, eviction causes irreversible information loss, potentially degrading generation quality.
KV Cache Merging. KV cache merging combines less important KV entries instead of permanently discarding them. DMC (nawrot2024dmc) learns when and how to merge through training, which limits generalization and introduces extra overhead. In contrast, CaM (zhang2024cam) adaptively merges evicted value states into others but does not merge the corresponding keys. Recently, D2O (wan2024D2O) selects merge candidates and assigns merging weights based on cosine similarity between key states, while KVMerger (wang2024KVMerger) introduces a clustering-based method to group merge candidates and computes merging weights using Gaussian Kernel Weights. However, these methods fail to maintain attention consistency before and after merging, leading to output perturbation. We propose a novel merging approach designed to eliminate output perturbation, supported by theoretical analysis and extensive comparisons.
@@ 3.1 Preliminary: Inference with KV Cache
$K_{t}$ $=[K_{t-1},k_{t}],\quad k_{t}=x_{t}W_{k}$
$V_{t}$ $=[V_{t-1},v_{t}],\quad v_{t}=x_{t}W_{v}$ (2)
$A^{t}$ $=\text{softmax}\left(\frac{q_{t}K_{t}^{T}}{\sqrt{d}}\right),\quad q_{t}=x_{t}W_{q}$
$s^{t}_{i}$ $=e^{\frac{q_{t}k_{i}}{\sqrt{d}}},\quad o_{t}=\sum_{i=1}^{t}A_{i}^{t}v_{i}=\frac{\sum_{i=1}^{t}s^{t}_{i}v_{i}}{\sum_{i=1}^{t}s^{t}_{i}}$ (3)
@@ 3.2 Rethinking KV Cache Eviction and Merging
Although current methods optimize eviction and cache allocation strategies (yang2024pyramidinfer; fu2024headkv) to minimize output impact, they cannot eliminate the perturbation in Equation 4. Previous studies have indicated that attention is not always sparse, especially in tasks requiring full context, as shown in Figure 2. Moreover, evicted KV may become important later, but irreversible eviction leads to permanent loss.
@@
Attention Sag in KV Cache Merging. Merging methods integrate less important KV into others rather than discarding them. Existing studies typically use weighted merging (nawrot2024dmc; wan2024D2O; wang2024KVMerger); formally, merging $(k_{e},v_{e})$ into $(k_{c},v_{c})$ is expressed as:
@@
Here, $(k_{r},v_{r})$ are the merged vectors, with weights $w_{e},w_{c}$ determined by the merging method. In D2O (wan2024D2O), they depend on the cosine similarity between $k_{e}$ and $k_{c}$, while in KVMerger (wang2024KVMerger), they are computed using Gaussian Kernel values. The weights satisfy the normalization condition $w_{e}+w_{c}=1$. However, this widely used convex combination method also introduces output perturbations:
Theorem 2.
@@ 3.3 Method: KeepKV
Electoral Votes and ZIP-Merging
Electoral Votes. To address Attention Sag, we propose the Electoral Votes mechanism, which records the number of merges $p_{i}$ (initialized to $1$) each KV pair undergoes. A natural analogy is the Electoral College system, where electors hold votes proportional to their state’s population rather than a uniform share. The attention score of each KV is then scaled by its votes to approximate the original multiple KV’s influence before merging. For example, if a KV pair $(k_{r},v_{r})$ has a vote count of $p_{r}=3$, it is equivalent to three identical and independent instances of $(k_{r},v_{r})$ participating in the attention computation. Formally, the outputs before ($o_{t}$) and after merging ($o^{\prime}_{t}$) are defined as follows:
$o_{t}$ $=\frac{\sum_{i=1}^{t}p_{i}s_{i}^{t}v_{i}}{\sum_{i=1}^{t}p_{i}s_{i}^{t}},$
$o^{\prime}_{t}$ $=\frac{\sum_{i=1,i\neq e,c}^{t}p_{i}s_{i}^{t}v_{i}+p_{r}s_{r}^{t}v_{r}}{\sum_{i=1,i\neq e,c}^{t}p_{i}s_{i}^{t}+p_{r}s_{r}^{t}},$
$p_{r}$ $=p_{e}+p_{c}.$ (6)
@@
$k_{r}$ $=\frac{(w_{e}k_{e}+w_{c}k_{c})\ln{\frac{w_{e}+w_{c}}{p_{e}+p_{c}}}}{w_{e}\ln{s_{e}^{t}}+w_{c}\ln{s_{c}^{t}}},$
$v_{r}$ $=\frac{w_{e}v_{e}+w_{c}v_{c}}{w_{e}+w_{c}},$
$w_{e}$ $=p_{e}s_{e}^{t},\quad w_{c}=p_{c}s_{c}^{t}.$ (7)
Theorem 3.
@@
This theorem confirms that our novel merging approach can eliminate output perturbations and complete the subtask introduced at the beginning of this section. However, its applicability remains limited to the current iteration, and extending it to multi-step generation requires additional design.
Extending to Multi-Step Generation
EMA Attention Scores. For ZIP-Merging to be effective in real-world multi-step generation, a solid comprehension of attention score dynamics is essential. Fortunately, empirical observations show that attention scores exhibit strong locality (Figure 2 (d)), meaning a token’s attention scores evolve smoothly across adjacent steps, which is also validated by prior studies (yang2024pyramidinfer; zhang2024cam; dong2024qaq). From this, we employ the Exponential Moving Average (EMA) (hunter1986EMA; busbridge2023EMAscale) with bias correction, a widely used technique in time-series analysis, formulated as follows:
[…]
Note that after the prefill stage, we compute EMA scores using a recent window of length $w$ rather than the entire sequence to obtain a more accurate estimation (li2024SnapKV; yang2024pyramidinfer). We find that this method outperforms mainstream approaches, such as cumulative attention and sliding window averaging, in predicting attention scores.Building on this, replacing all score $s_{i}^{t}$ in Equation 7 with our EMA scores $\hat{s_{i}^{t}}$ from Equation 8 successfully achieves the extension to multi-step generation. Consequently, the future output perturbation becomes estimable and controllable. We present the following theorem and lemma(proof in Appendix A.4):
Theorem 5.
@@
Algorithm 1 KeepKV Merging at $t$-th Step
1:Input: Attention scores $s^{t}$, EMA scores $S^{t-1}$, KV cache
2: Let $K_{e}$ denote the to-be-evicted cache
3: Let $K_{c}$ denote the retained cache
[…]
5: For each $k_{e}\in K_{e}$, select $k_{c}=\operatorname*{Argmax}_{k_{c}\in K_{c}}(U^{e}),\ U_{e,c}>T$
[…]
7:Merge:
8:  $k_{r}=\frac{\ln{((p_{e}\hat{s_{e}^{t}}+p_{c}\hat{s_{c}^{t}})/(p_{e}+p_{c}))}}{p_{e}\hat{s_{e}^{t}}\ln{\hat{s_{e}^{t}}}+p_{c}\hat{s_{c}^{t}}\ln{\hat{s_{c}^{t}}}}(p_{e}\hat{s_{e}^{t}}k_{e}+p_{c}\hat{s_{c}^{t}}k_{c})$
9:  $v_{r}=\frac{1}{p_{e}\hat{s_{e}^{t}}+p_{c}\hat{s_{c}^{t}}}(p_{e}\hat{s_{e}^{t}}v_{e}+p_{c}\hat{s_{c}^{t}}v_{c})$
10:  $p_{r}=p_{e}+p_{c}$
11: Discard $(k_{e},v_{e}),(k_{c},v_{c})$ and insert $(k_{r},v_{r})$ into KV cache
12:Output: Updated KV cache
@@
Similarity-driven merging. Lemma 6 shows that output perturbation decreases as prediction error $\epsilon$ reduces, and closer merging objects result in lower perturbation. Clearly, if the merged KV pairs are identical, retaining one pair and setting its Electoral Votes to 2 introduces no error in subsequent computations. This provides a theoretical justification for prior merging strategies favoring high-similarity KV pairs (wan2024D2O; wang2024KVMerger). Following this, we merge each evicted KV pair with the retained one having the highest cosine similarity of keys, using a predefined threshold $T$ to determine whether merging should occur, avoiding the overhead of dynamic adjustments like D2O (wan2024D2O). Furthermore, we observe that, during the prefill stage, reversing the conventional order—by first merging based on key similarity and then applying the eviction strategy, instead of merging after eviction as commonly done—can improve generation quality.
We present the workflow of KeepKV in Figure 3. Notably, KeepKV imposes no specific constraints on cache allocation or token selection strategies. It can directly integrate with common token selection methods by designating the merging pairs based on their eviction and retention sets, and it is also compatible with various cache allocation strategies. Thus, KeepKV demonstrates strong adaptability and can be combined with a range of mainstream cache compression methods, significantly enhancing both compression capability and generation quality.